# Distilling the Implicit Multi-Branch Structure in LLMs’ Reasoning via Reinforcement Learning

### v1

@@ Appendix A Ablation study of Structured Fine-grained Training
Figure 6 shows the ablation study of our Structured Fine-grained Training method by showing the accuracy and structure matching varying with training steps. It indicates that generative structure reward model trained by our structured fine-grained training method can effectively optimize the task accuracy (Figure 6) and reasoning structure matching (Figure 6) between the teacher and the student a, which grows progressively as the training steps.
(a) Accuracy
@@ Appendix B Ablation study for Reward Weights
In this section, we conduct ablation study about the reward weights in RL training. The essential rewards in our method include accuracy reward ($R_{acc}$) for the specific tasks such as rule-based accuracy reward in math, and our structure matching reward ($R_{gsrm}$) obtained from generative structure reward model, so we primarily discuss the weight relationship between these two rewards. Specifically, we adjust the proportional relationship between $R_{gsrm}$ and $R_{acc}$ ($\frac{R_{gsrm}}{R_{acc}}$) from $0.2$ to $2$ in increments of $0.2$, and observe the variations in pass@1 (n=16) and pass@16 (n=16) on AIME24. This experimental results are shown in Figure 7. We can see that when the weights of the two are equal $\frac{R_{gsrm}}{R_{acc}}=1$, the LLM achieves a harmonized optimal effect in both the correctness of reasoning (pass@1)and the capability of sampling (pass@16). So in our experiments, we set the reward weights as: $R_{acc}:R_{gsrm}:R_{format}:R_{tag}=3:3:2:2$, in which $R_{format}$ is format reward and $R_{tag}$ is tag count reward that assesses whether the thought content adheres to the specified format Guo et al. [2025].

### v2

@@ Appendix A Ablation study of Structured Fine-grained Training
Figure 6 shows the ablation study of our Structured Fine-grained Training method by showing the accuracy and structure matching varying with training steps. It indicates that generative structure reward model trained by our structured fine-grained training method can effectively optimize the task accuracy (Figure 6) and the matching degree of reasoning structure (Figure 6) between the teacher and the student, which grows progressively as the training steps.
(a) Combined Accuracy (result acc. + format acc.)
@@ Appendix B Ablation study for Reward Weights
In this section, we conduct ablation study about the reward weights in RL training. The essential rewards in our method include accuracy reward ($R_{acc}$) for the specific tasks such as rule-based accuracy reward in math, and our structure matching reward ($R_{gsrm}$) obtained from generative structure reward model, so we primarily discuss the weight relationship between these two rewards. Specifically, we adjust the weight between $R_{gsrm}$ and $R_{acc}$ ($\frac{R_{gsrm}}{R_{acc}}$) from $0.2$ to $2$ in increments of $0.2$, and observe the variations in pass@1 (n=16) and pass@16 (n=16) on AIME24. This experimental results are shown in Figure 7. We can see that when the weights of the two are equal, i.e., $\frac{R_{gsrm}}{R_{acc}}=1$, the LLM achieves a harmonized optimal effect in both the correctness of reasoning (pass@1)and the capability of sampling (pass@16). So in our experiments, we set the reward weights as: $R_{acc}:R_{gsrm}:R_{format}:R_{tag}=3:3:2:2$, in which $R_{format}$ is format reward and $R_{tag}$ is tag count reward that assesses whether the thought content adheres to the specified format Guo et al. [2025].